OpenStudy (anonymous):

What is the exact value of the expression 7√8 + 2 √12 - 3 √2? Show your work

OpenStudy (anonymous):

\[7\sqrt{8}+2\sqrt{12}-3\sqrt{2}\] \[7\sqrt{4}\sqrt{2}+2\sqrt{4}\sqrt{3}\] \[14\sqrt{2}+4\sqrt{3}-3\sqrt{2}\] \[11\sqrt{2}+4\sqrt{3}\] If this is not it then i dont know because i simplified it as best as i could by taking out square roots

OpenStudy (anonymous):

Thanks alot

OpenStudy (anonymous):

Did it help?

OpenStudy (anonymous):

Alot

OpenStudy (anonymous):

good! :)